How many cubic centimeters can a cigar box hold if its dimensions are 3 centimeters by 13 centimeters by 15 centimeters? Include correct units with your solution. Show work.
Answer by checkley77(12844)

V=L*W*H
V=3*13*15=1,170 CM^3. ANS.